Session: May 2003 (session 2)
Wavelength: 6 cm
To give feedback on this experiment, click
Medicina (andrea at Tue Jun 3 01:30:52 2003) Minor failures :
Shanghai (Xinyong Huang at Tue Jun 3 02:36:20 2003) Failure :
The experiment was not observed at Shanghai station, because of the antenna with trouble.
Torun (K.Borkowski at Tue Jun 3 04:33:03 2003) Success :
No known problems.
Onsala (Michael Lindqvist at Tue Jun 3 06:08:06 2003) Minor failures : RECO
Tape rolled of at 21.17 UT, missed some scans, see log.
Medicina (G.Maccaferri at Tue Jun 3 07:53:50 2003) Minor failures : RECO
At 20:11, the tape change, the schedule commanded a "post-pass" (it takes about 22min) allowing only 15 minutes for the whole procedure, tape changed included. This casused the lost of first 3 scans of tape #2, and tape unthread at end of first pass.
In future, please reserve more time in case of postpass
Urumqi (Liu Xiang at Tue Jun 3 08:00:37 2003) Minor failures : TSYS HIGH PCAL FS
no p_cal. and the cooling system seems not working properly, so Tsys
is about 140-160K measured with radiometer; however, FS gives wrong
numbers of Tsys in BBCs because some problem in FS, so you could
use Tsys in if1 and if2 in FS and use a factor for correction.
Effelsberg (A. Kraus at Tue Jun 3 09:14:42 2003) Minor failures : RECE FS
Data between UT1258-1324 was lost due to receiver problems. Afterwards the FS crashed 5 times (roughly every 2 hours) without clear reason (the problem is under investigation). Each time we lost about 15 minutes.
Noto (P. Leto at Tue Jun 3 11:11:51 2003) Success :
Lost from 20:28:45 to 20:51:15 of DOY 153 there was not enough time
for tape changing.
Urumqi (Liu Xiang at Tue Jun 3 13:17:51 2003) Minor failures : TSYS FS
Comment: the FS has some problem with Tsys measure, the values follow i1 and i2 are roughly correct; the values we measured with radiometer are around 140-160K, so you could use a factor from this in ANTAB.
Jodrell_Bank (Alastair Gunn at Wed Jun 4 23:55:06 2003) Minor failures : RECO
Formatter problem between 15:40 and 16:00 UT. Many problems
with vacuum failure. Failed at 17:32, restarted at 18:35 UT.
Also failed at 20:02 and 21:20 UT.
Westerbork (foley at Mon Jul 7 15:39:24 2003) Minor failures : FS
FS forced a postpass during tape change. Lost 20:11 to 21:36
with re-synchinhg the tape